# Imaging heat transport in suspended diamond nanostructures with integrated spin defect thermometers

### v1

@@ Abstract
Among all materials, mono-crystalline diamond has one of the highest measured thermal conductivities, with values above 2000 W/m/K at room temperature. This stems from momentum-conserving ‘normal’ phonon-phonon scattering processes dominating over momentum-dissipating ‘umklapp’ processes, a feature that also suggests diamond as an ideal platform to experimentally investigate phonon heat transport phenomena that violate Fourier’s law. Here, we introduce dilute nitrogen-vacancy colour centers as in-situ, highly precise spin defect thermometers to image temperature inhomogeneities in single-crystal diamond microstructures heated from ambient conditions. We analyse cantilevers with cross-sections in the range from about 0.2 to 2.6 m2, observing a relation between cross-section and heat flux that departs from Fourier’s law predictions. We rationalize such behavior relying on first-principles simulations based on the linearized phonon Boltzmann transport equation, also discussing how fabrication-induced impurities influence conduction. Our temperature-imaging method can be applied to diamond devices of arbitrary geometry, paving the way for the exploration of unconventional, non-diffusive heat transport phenomena.
@@
Carbon-based materials such as graphene, graphite or diamond present exceptional heat-conduction properties. Due to large phonon group velocities, they exhibit high thermal conductivity – with diamond reaching record values at room temperature and below [1, 2, 3]. Moreover, the prevalence of normal (momentum-conserving) phonon-phonon scattering processes over Umklapp (momentum non-conserving) processes in these materials leads to spectacular phonon heat transport phenomena, that violate Fourier’s law for macroscopic diffusive heat propagation [4]. In graphene and graphite, ballistic transport in nanoscale devices [5, 6] or hydrodynamic heat transport [7, 8, 9, 10, 11] have been evidenced at or near room temperature. While recent theoretical work predicted that diamond should also exhibit heat transport deviating from Fourier’s diffusion at room temperature [12], little experimental work has been reported in this direction so far. Noteworthy, the influence of phonon scattering at grain boundaries has been explored in polycrystalline diamond [13] and unexpected size effects were recently reported in the thermal conductivity of diamond microparticles [14]. Understanding heat transport in nanostructured diamond is also valuable in view of its applications in optomechanics [15, 16], nonlinear optics [17], quantum nanophotonics [18], as well as in heat management devices [19].
@@
A major challenge in the study of thermal transport at the nanoscale is to find appropriate temperature sensors. Traditional measurements of the heat conductivity rely on resistive sensors combined with advanced sample engineering [5, 9] and cannot provide spatially-resolved temperature measurements. Time-resolved non contact optical thermometry techniques such as transient thermal gratings [20, 7], and time- or frequency domain thermoreflectance [21, 22, 23] have excellent temporal resolution but limited spatial resolution, and are often applicable to simple planar geometries only. Other optical techniques [24, 25], including Raman sideband thermometry [26, 27, 28], offer good spatial resolution but rather poor temperature accuracy, typically not better than a few K.
@@
Fittingly, diamond hosts optically active spin defects, the most studied being the negatively charged nitrogen vacancy (NV) center [29], which can be used as in-situ nanoscale thermometers [30, 31, 32]. Negatively charged NV centers are spin-1 atomic defects formed by a substitutional nitrogen atom neighboring a vacancy in the diamond lattice. Their spin can be initialized and readout by optical pumping and monitoring the photoluminescence intensity. The resonance frequency associated with the transition between and spin states is temperature-dependent [33], as can be tracked with either continuous-wave or pulsed optically detected magnetic resonance (ODMR) [34]. NV centers are thus excellent thermometers, offering nanoscale spatial resolution and sub- sensitivity [35] for steady-state temperature measurement near room temperature. Solid-state spin defects acting as temperature sensors are also found in a broad range of materials, including hBN [36], Si [37], SiC [38, 39, 40] and this field is currently undergoing rapid developments. However, the potential of quantum sensors in the study of non-Fourier thermal transport remains largely unexplored and unexploited.
@@
In this work, we use dilute NV centers in the diamond lattice as precise in-situ temperature sensors to investigate heat transport properties of suspended diamond microstructures. The principle of our experiment is summarized in Fig.1a. We fabricate suspended micro-cantilevers out of single crystal diamond grown by chemical vapour deposition and enriched in (%), containing NV centers up to a concentration of 3 ppm. Cantilevers are patterned using electron-beam lithography and subsequent etching, in two steps: first a step of vertical, oxygen-based reactive ion etching, followed by focused ion beam (FIB) milling at an angle from the diamond surface. This results in triangular cantilever cross-sections, as visible in Fig. 1b. The cantilever acts as a one-dimensional channel for heat transport. A small chromium patch is added at the cantilever tip to serve as a heat source, activated by optical absorption of a 515 nm laser focused onto it, referred to as heating laser. The bulk diamond acts as a heat sink. Complete details on the sample fabrication can be found in the Supplementary Material (SM) [41].
@@
Photoluminescence from NV centers is collected in confocal geometry, using a second 515 nm laser beam to address a large ensemble of NV centers (about ) located in the diffraction-limited spot. This second laser is referred to as readout laser. The temperature-dependent zero-field splitting between and states is measured through continuous wave (cw) ODMR. We use a lock-in amplifier (LIA) combined with frequency modulation of the microwave to achieve high signal to noise ratio [42]. Fig. 1c presents the shift of ZFS value versus temperature, defined as , measured both in bulk diamond and on cantilevers. Note that, at room temperature, the ZFS value is around 2.87 GHz. A position-dependent offset can be observed, due for example to the presence of local strain, particularly in cantilevers. As visible in Fig. 1c, this offset does not modify the temperature-dependent relative ZFS shift. Our data is matching previous reports, and the ZFS shift is fitted by the model from Ref. [43], with only two fitting parameters (dashed line in Fig. 1c).
@@
The calibration is used to convert values of ZFS shift to local temperature under the readout laser spot. Noise analysis gives a typical temperature sensitivity of 0.27 for the temperature readout on the cantilever [41]. Note that the temperature sensitivity depends on the number of NV probed by the readout spot, it can thus degrade when addressing smaller cantilevers, but also vary with the NV density in the starting diamond substrate. The choice of cw-ODMR is a trade-off between sensitivity requirement and simplicity of experimental implementation. Even though more sophisticated NV-based thermometry schemes allow to reach better sensitivity, relying e.g. on pulse sequences [30, 31] or on multi-frequency drive [35], in our case uncertainties in the absorbed laser power and exact cantilever cross-sections have larger impacts on the final error bars, as detailed below.
@@
The heating laser is then used to induce a heat flow in a cantilever. Fig. 2a presents the measured local temperature shift as a function of heating laser power , for a readout spot located at a fixed position, in the middle of the cantilever. We observe a linear relation between and heating power. Standard deviations of residuals after a linear fit give a temperature error of K [41]. This value includes systematic errors due e.g to imperfect alignment of the heating laser on the gold patch, explaining why it is slightly higher than the readout error ( mK with our 30 s integration time).
@@
Fixing the heating laser power, we measure the space-dependent temperature profile along the cantilever by sweeping the position of the readout laser. For each readout laser position, two measurements are necessary to extract the local heating-induced temperature shift: is obtained by subtracting the ZFS value without and with heating. The temperature profile for a cantilever that has a narrower section in the middle, see Fig. 2b, is shown in Fig. 2c. A section of constant width shows linear temperature profile. Moreover, the narrow section has a much steeper temperature gradient than the wide section, while the temperature is constant on the free-hanging side, away from the Cr patch. In our experimental setting, the spatial resolution of the temperature measurement is diffraction-limited, the readout laser spot size being 0.44 µm.
We emphasize that the ZFS can also shift due to local strain or stress. This is for example visible in the evolution of ZFS along the cantilever (in the absence of heating), where a strong variation is observed near the cantilever base [41]. However, since our cantilevers are free-standing, we do not expect that heating induces any additional strain except near its base. We also confirmed that the Cr patch is well thermalized with the diamond by independent estimation of its temperature using anti-Stokes emission spectroscopy [41].
@@
We can consider the case of standard diffusive heat transport to develop a picture of the behavior expected from the established Fourier’s law. According to Fourier’s law, in 1D, , with the local heat flux density and the material’s thermal conductivity. Neglecting all other thermal losses (mainly convective or radiative), Fourier’s law predicts a constant heat flux along the cantilever , with the heating laser power incident on the Cr patch, the Cr patch absorption coefficient and the cantilever cross-section. is temperature dependent, but for small heating power, Fourier’s law predicts a linear temperature profile along the cantilever, with a local temperature shift directly proportional to the heating laser power, compatible with the results from Fig. 2. However, having a linear temperature profile is a necessary but not sufficient condition for the validity of Fourier’s law. In fact, Fourier’s law also assumes thermal conductivity to be an intrinsic, size-independent material’s property, and consequently a heat flux that scales linearly with the cantilever cross-section. On the contrary, when reducing the lateral size of a 1D channel, we intuitively expect size effect to become increasingly more relevant, leading to a reduction of . In the following we investigate the heat flux’s magnitude as a function of cantilever width.
@@
In diamond, has been predicted to decreases to approximately half of its bulk value at widths 0.7m [44, 45] – a much stronger reduction than other materials for this channel size[46, 47, 48, 49, 50]. However, the width dependence of thermal conductivity in diamond has not been experimentally demonstrated yet. Our experimental platform is ideally suited for investigating this phenomenon. We investigate cantilevers with different lateral width , in the range 0.8-2.5 µm, and identical length of 10 µm. The lower bound of this range originates from limitations in temperature sensitivity for cantilever widths below the diffraction limited spot size (due to the reduced number of probed NV centers). Conversely, the upper bound is due to difficulty in fabricating suspended cantilevers with dimensions bigger than 3 µm (due to the long etching time required). We measure the temperature profile along each individual cantilever, adjusting such that the total temperature shift is comparable in all cases and remains below 40 K. The results are presented in Fig. 3, with panel c presenting the measured temperature gradient , normalized by , versus cantilever width . In the case of purely diffusive thermal transport, Fourier law predicts that this normalized temperature gradient is inversely proportional to the cantilevers cross section. Specifically, since all our cantilevers have triangular cross section with same angle , Fourier’s law predicts . In contrast, our experiments deviate from Fourier’s prediction, with a temperature gradient displaying a scaling approximatively proportional to for µm. This indicates that the probed range of cantilever width is within the non-Fourier regime, where non-diffusive effects are relevant and thermal conductivity depends on the sample’s geometry (size).
@@
In Fig. 4 we compare the experimental values for , extracted from the measurements in Fig. 3 (see SM [41] for details), against predictions obtained solving the VHE or Fourier’s law. In the numerical simulations, we consider two different impurity scenarios to compute transport coefficients from first principles. First, we consider the ideal case of pristine diamond cantilevers with only 0.05% 13C isotopes (dashed lines). In this idealized case, VHE and Fourier’s law yield values having practically unimportant differences, however both significantly higher than the experimental values.
@@
To address this discrepancy, we investigate theoretically the effect of possible additional sources of thermal resistance missing from our simulations in pristine samples. In particular, the cantilever’s fabrication procedure based on FIB milling is known to cause contamination of Ga ions in proximity of the milled surfaces [57] (Fig. 4 inset). Ga contamination can be as high as hundreds of ppm within a penetration layer of depth estimated in the present case to µm [41]. The penetration layer is determined by the FIB specifications, it is therefore the same for cantilevers having different cross sections. This implies that the average concentration of Ga impurities in our samples is proportional to the surface-to-volume ratio. Consequently, Ga contamination is expected to have an increasingly more important effect as the cantilevers width decreases. Therefore, we account for the influence of Ga impurities using the model by Ref. [56], considering an average concentration of Ga impurities proportional to the surface-to-volume ratio, and determine one parameter — the concentration of Ga impurities in the penetration layer — from a fit with experimental data [41]. NV center impurities also present in our diamond samples are negelected due to their small concentration (3 ppm) [58].
@@
Solid lines in Fig. 4 are computed values for obtained from this one-parameter model, where the concentration of Ga impurities in the penetration layer obtained from the fit is ppm (a value of the same order of magnitude of those discussed in Ref. [57]). We highlight the good quantitative agreement between predictions and experiments. Contrarily, a model with constant concentration of impurities over the whole cross-section fails to reproduce the experiments, especially in the trend [41]. This suggests that the main hypothesis of our model — having average concentration of Ga impurities proportional to the surface-to-volume ratio of the cantilevers — is realistic and captures the most important microscopic physics. Finally, hydrodynamic effects (quantified by the difference between green and red curves) appear practically unimportant and the deviation from Fourier’s law are mainly caused by being not an intrinsic property but a size-dependent quantity.
@@
In conclusion, we highlight that this platform offers multiple routes for new investigations. First, the question arises of how to maximize heat conductivity in sub-micrometer diamond channels, which can be answered by achieving a finer control on the impurity level. The effect of sidewall roughness on phonon scattering at the boundaries [59, 60] can also be probed by varying the etching method and chemical termination of the diamond surface. Second, hydrodynamic effects in heat transport could be revealed using refined geometries that enhance unambiguous signatures of hydrodynamic features, such as heat vortices [54]. Third, this research can be pushed toward even smaller channels to address open questions in quantized phononic heat conductance [61]. Overall, our results demonstrate the potential of solid state spin defects in ultrahigh thermal conductivity materials to explore heat transport phenomena in non-Fourier regimes.
@@ Acknowledgements.
Acknowledgments. This project has received funding from the Swiss National Science Foundation through grant #198898 and #216406, EPFL Center for Quantum Science and Engineering through a Collaborative Research Fellowship, the European Union’s Horizon 2020 research and innovation programmes under the Marie Skłodowska-Curie grant agreement #945363 and under grant agreement #101112347, Project NerveRepack. M. S. acknowledges support from: (i) the Kelvin2 HPC platform at the NI-HPC Centre (funded by EPSRC and jointly managed by Queen’s University Belfast and Ulster University); (ii) the UK National Supercomputing Service ARCHER2, for which access was obtained via the UKCP consortium and funded by EPSRC [EP/X035891/1]; (iii) Gonville and Caius College. N.M and E.D. acknowledge support from the Swiss National Science Foundation (SNSF), through Grant No. CRSII5_189924 (“Hydronics” project). The diamond cantilevers were fabricated at the EPFL Center of MicroNanoTechnology (CMi) and we gratefully acknowledge the support of EPFL CMi.References
@@
- [41] See Supplemental Material at … for further details on experimental and theoretical methods and additional data, which includes Refs. [62-82].

### v2

@@ Abstract
Among all materials, mono-crystalline diamond has one of the highest measured thermal conductivities, with values above 2000 W/m/K at room temperature. This stems from momentum-conserving ‘normal’ phonon-phonon scattering processes dominating over momentum-dissipating ‘Umklapp’ processes, a feature that also suggests diamond as an ideal platform to experimentally investigate phonon heat transport phenomena that violate Fourier’s law. Here, we introduce dilute nitrogen-vacancy color centers as in-situ, highly precise spin defect thermometers to image temperature inhomogeneities in single-crystal diamond microstructures heated from ambient conditions. We analyze cantilevers with cross-sections in the range from about 0.2 to 2.6 µm2, observing a relation between cross-section and heat flux that departs from Fourier’s law predictions. We rationalize such behavior relying on first-principles simulations based on the linearized phonon Boltzmann transport equation, also discussing how fabrication-induced impurities influence conduction. Our temperature-imaging method can be applied to diamond devices of arbitrary geometry, paving the way for the exploration of unconventional, non-diffusive heat transport phenomena.
@@
Carbon-based materials such as graphene, graphite or diamond present exceptional heat-conduction properties. Due to large phonon group velocities, they exhibit high thermal conductivity – with diamond reaching record values at room temperature and below [1, 2, 3]. Moreover, the prevalence of normal (momentum-conserving) phonon-phonon scattering processes over Umklapp (momentum non-conserving) processes in these materials leads to phonon heat transport phenomena that violate Fourier’s law for macroscopic diffusive heat propagation [4]. In graphene and graphite, ballistic transport in nanoscale devices [5, 6] or hydrodynamic heat transport [7, 8, 9, 10, 11, 12, 13] have been evidenced at or near room temperature, and recent theoretical work predicted that diamond should also exhibit hydrodynamic heat transport at room temperature [14]. The influence of phonon scattering at grain boundaries has been explored in polycrystalline diamond [15] and unexpected size effects were recently reported in the thermal conductivity of diamond microparticles [16], but experimental data on low-dimensional single-crystal diamond is lacking. Understanding heat transport in nanostructured diamond will be valuable in view of its applications in optomechanics [17, 18], nonlinear optics [19], quantum nanophotonics [20], as well as in heat management devices [21].
@@
A major challenge in the study of thermal transport at the nanoscale is to find appropriate temperature sensors. Traditional measurements of the heat conductivity rely on resistive sensors combined with advanced sample engineering [5, 11] and cannot provide spatially-resolved temperature measurements. Non contact optical thermometry techniques such as transient thermal gratings [22, 9], and time- or frequency-domain thermoreflectance [23, 24, 25] have been widely employed on simple planar geometries but require the deposition of a transducer material, typically a metal [26], and don’t provide a direct temperature readout [27]. Other optical techniques [28, 29], including Raman sideband thermometry [30, 31, 32, 33], offer good spatial resolution but rather poor temperature accuracy, typically not better than a few K.
@@
In this work, we use dilute color centers in the diamond lattice as precise in-situ temperature sensors to investigate heat transport properties of suspended diamond microstructures. Diamond hosts optically active spin defects, the most studied being the negatively charged nitrogen vacancy (NV) center [34], which can be used as in-situ nanoscale thermometers [35, 36, 37]. Negatively charged NV centers are spin-1 atomic defects formed by a substitutional nitrogen atom neighboring a vacancy in the diamond lattice. Their spin can be initialized and readout by optical pumping and monitoring the photoluminescence intensity. The resonance frequency associated with the transition between and spin states is temperature-dependent [38], as can be tracked with either continuous-wave or pulsed optically detected magnetic resonance (ODMR) [39]. NV centers are thus excellent thermometers, offering nanoscale spatial resolution and sub- sensitivity [40] for steady-state temperature measurement near room temperature. Solid-state spin defects acting as temperature sensors are also found in a broad range of materials, including hBN [41], Si [42] and SiC [43, 44, 45], making this method widely applicable.
@@
The principle of our experiment is summarized in Fig.1a. We fabricate suspended micro-cantilevers out of single crystal diamond grown by chemical vapour deposition and enriched in (%), containing NV centers up to a concentration of 3 ppm. Cantilevers are patterned using electron-beam lithography and subsequent etching, in two steps: first a step of vertical, oxygen-based reactive ion etching, followed by focused ion beam (FIB) milling at an angle from the diamond surface. This results in triangular cantilever cross-sections, as visible in Fig. 1b. The cantilever acts as a one-dimensional channel for heat transport. A small chromium patch is added at the cantilever tip to serve as a heat source upon absorption of a 515 nm laser focused onto it (referred to as heating laser) and the bulk diamond acts as a heat sink. In the future, a UV-C laser could be used to heat diamond directly, as shown in other wide-gap materials [33]. Complete details on the sample fabrication can be found in the Supplementary Material (SM) [46].
@@
Photoluminescence from NV centers is collected in confocal geometry, using a second 515 nm laser beam to address a large ensemble of NV centers (about ) located in the diffraction-limited spot. This second laser is referred to as readout laser. The temperature-dependent zero-field splitting (ZFS) between and states is measured through continuous wave (cw) ODMR. We use a lock-in amplifier (LIA) combined with frequency modulation of the microwave to achieve high signal to noise ratio [47]. The absolute value of the ZFS is found to depend on position [46], possibly due to the presence of local strain. Since our cantilevers are free-standing, we do not expect that heating induces any additional strain (except maybe very near its base). We measure the change of ZFS against a reference temperature and present in Fig. 1c the results for bulk diamond and a cantilever. Our data is fitted by the model from Ref. [48], with only two fitting parameters (dashed line in Fig. 1c).
@@
The calibration is used to convert values of ZFS shift to local temperature under the readout laser spot. Noise analysis gives a typical temperature sensitivity of 0.27 for the temperature readout on the cantilever [46]. The choice of cw-ODMR is a trade-off between sensitivity requirement and simplicity of experimental implementation. Even though more sophisticated NV-based thermometry schemes allow to reach better sensitivity, relying e.g. on pulse sequences [35, 36] or on multi-frequency drive [40], the dominant sources of uncertainties in our case come from the estimates of absorbed laser power and cantilever cross-section, as detailed below.
@@
The heating laser is used to induce a heat flow in a cantilever. Fig. 2a presents the measured local temperature rise as a function of heating laser power , for a readout spot located at a fixed position in the middle of the cantilever. We observe a linear relation between and heating power. Standard deviations of residuals after a linear fit give a temperature error of K [46]. This value includes systematic errors due e.g to imperfect alignment of the heating laser on the gold patch, explaining why it is slightly higher than the readout error ( K with our 3 s integration time).
@@
Fixing the heating laser power, we measure the space-dependent temperature profile along the cantilever by sweeping the position of the readout laser. For each readout laser position, the ZFS values are measured without and with heating. To illustrate the method, the temperature profile for a cantilever of varying width is shown in Fig. 2b-c. A section of constant width shows linear temperature profile and the narrow section has a much steeper temperature gradient than the wide section, while the temperature is constant on the free-hanging side. In our experimental setting, the spatial resolution is diffraction-limited, the readout laser spot size being 0.44 µm in diameter.
@@
We can consider the case of standard diffusive heat transport to develop a picture of the expected behavior: According to Fourier’s law, in 1D, , with the local heat flux density and the material’s thermal conductivity, which depends slowly on temperature. We confirmed that the Cr patch’s temperature remains close to that of the underlying diamond [46], so that convective and radiative thermal losses can be neglected. In this case, , with the heating laser power incident on the Cr patch, the Cr patch absorption coefficient and the cantilever cross-section. For small , Fourier’s law predicts a linear temperature profile along the cantilever, compatible with the results from Fig. 2. However, having a linear temperature profile is a necessary but not sufficient condition for the validity of Fourier’s law, since it also assumes thermal conductivity to be an intrinsic, size-independent material’s property, and consequently predicts a heat flux that scales linearly with the cantilever cross-section. Instead, in diamond, has been predicted to decrease to approximately half of its bulk value at widths 0.7m [49, 50] – a much stronger reduction than other materials for this channel size[51, 52, 53, 54, 55]. We now show that a reduction of is observed when reducing the lateral size of the cantilever.
@@
We investigate cantilevers with different lateral width , in the range 0.8-2.5 µm, and identical length of 10 µm. The lower bound of this range originates from limitations in temperature sensitivity for cantilever widths below the diffraction limited spot size (due to the reduced number of probed NV centers). Conversely, the upper bound is due to the difficulty in fabricating suspended cantilevers with dimensions bigger than 3 µm (due to the long etching time required). We measure the temperature profile along each individual cantilever, adjusting such that the maximum temperature rise is comparable in all cases and remains below 40 K. The results are presented in Fig. 3, with panel (c) presenting the measured temperature gradient , normalized by , versus cantilever width . In the case of purely diffusive thermal transport, Fourier law predicts that this normalized temperature gradient is inversely proportional to the cantilevers cross-section. Specifically, since all our cantilevers have triangular cross-section with same angle , Fourier’s law predicts . In contrast, our experiments deviate from this prediction, with a temperature gradient displaying a scaling approximately proportional to for µm.
@@
In Fig. 4 we compare the experimental values for , extracted from the measurements in Fig. 3 (see SM [46] for details), against predictions obtained solving the VHE or Fourier’s law. In the numerical simulations, we consider two different impurity scenarios to compute transport coefficients from first principles. First, we consider the ideal case of pristine diamond cantilevers with only 0.05% 13C isotopes (dashed lines). In this idealized case, the VHE and Fourier’s law yield values having minor differences; however, both are significantly higher than the experimental values.
@@
To address this discrepancy, we investigate theoretically the effect of additional sources of thermal resistance missing from simulations of pristine samples. In particular, FIB milling is known to cause contamination by Ga ions [62] (Fig. 4 inset) that can be as high as hundreds of ppm within a penetration depth that we estimate to be µm [46]. This penetration layer is expected to be the same for all cantilevers since we used the same FIB parameters. Consequently, Ga contamination is expected to have an increasing impact on thermal conductivity as the cantilever width decreases. We account for the influence of Ga impurities using the model of Ref. [61], considering an average concentration of Ga impurities proportional to the surface-to-volume ratio, and determine a single parameter — the concentration of Ga impurities in the penetration layer — from a fit to the experimental data [46].
@@
Solid lines in Fig. 4 are the calculated values for within this one-parameter model, where the concentration of Ga impurities in the penetration layer obtained from the fit is ppm (a value of the same order of magnitude as those in Ref. [62]). Nitrogen and vacancy impurities are also present in the diamond with concentrations of the order of 10 ppm, which is expected to have a comparatively negligible effect on thermal conductivity [63]. A model with constant concentration of impurities over the whole cross-section fails to reproduce the experiments, especially in the steepness of the reduction of with decreasing [46]. This suggests that the main hypothesis of our model — having average concentration of Ga impurities proportional to the surface-to-volume ratio of the cantilevers — is realistic and captures the most important microscopic physics. Finally, hydrodynamic effects (quantified by the difference between green and red curves) appear practically unimportant and the deviation from Fourier’s law are mainly caused by being not an intrinsic property but a size-dependent quantity.
@@
In conclusion, we highlight that this experimental platform consisting of NV centers in diamond nanostructures offers multiple routes for new investigations. First, the question arises of how to maximize heat conductivity in sub-micrometer diamond channels, which can be answered by achieving a finer control on the impurity levels. The effect of sidewall roughness on phonon scattering at the boundaries [64, 65] can also be probed by varying the etching method and chemical termination of the diamond surface. Second, hydrodynamic effects in heat transport could be revealed using refined geometries that enhance unambiguous signatures of hydrodynamic features, such as heat vortices [59]. Third, this research can be pushed toward even smaller channels to address open questions in quantized phononic heat conductance [66]. Overall, our results demonstrate the potential of solid state spin defects in ultrahigh thermal conductivity materials to explore heat transport phenomena in non-Fourier regimes.
@@ Acknowledgements.
Acknowledgments. This project has received funding from the Swiss National Science Foundation through grant 198898 and 216406, EPFL Center for Quantum Science and Engineering through a Collaborative Research Fellowship, the European Union’s Horizon 2020 research and innovation programmes under the Marie Skłodowska-Curie grant agreement 945363 and under grant agreement 101112347, Project NerveRepack. M. S. acknowledges support from: (i) the Kelvin2 HPC platform at the NI-HPC Centre (funded by EPSRC and jointly managed by Queen’s University Belfast and Ulster University); (ii) the UK National Supercomputing Service ARCHER2, for which access was obtained via the UKCP consortium and funded by EPSRC [EP/X035891/1]; (iii) Gonville and Caius College. N.M and E.D. acknowledge support from the Swiss National Science Foundation (SNSF), through Grant No. CRSII5_189924 (“Hydronics” project). The diamond cantilevers were fabricated at the EPFL Center of MicroNanoTechnology (CMi) and we gratefully acknowledge the support of EPFL CMi.References
@@
- Lee et al. [2015] S. Lee, D. Broido, K. Esfarjani, and G. Chen, Hydrodynamic phonon transport in suspended graphene, Nature Communications 6, 6290 (2015).
@@
- Wilson et al. [2012] R. Wilson, B. A. Apgar, L. W. Martin, and D. G. Cahill, Thermoreflectance of metal transducers for optical pump-probe studies of thermal properties, Optics express 20, 28829 (2012).
- Huang et al. [2024] X. Huang, R. Anufriev, L. Jalabert, K. Watanabe, T. Taniguchi, Y. Guo, Y. Ni, S. Volz, and M. Nomura, A graphite thermal tesla valve driven by hydrodynamic phonon transport, Nature , 1 (2024).
@@
- Elhajhasan et al. [2023] M. Elhajhasan, W. Seemann, K. Dudde, D. Vaske, G. Callsen, I. Rousseau, T. F. Weatherley, J.-F. Carlin, R. Butté, N. Grandjean, et al., Optical and thermal characterization of a group-iii nitride semiconductor membrane by microphotoluminescence spectroscopy and raman thermometry, Physical Review B 108, 235313 (2023).
@@
- [46] See Supplemental Material at … for further details on experimental and theoretical methods and additional data.